OpenStudy (anonymous):

35 hours per week, one and a quarter hours per day, five days a week. How many days will it take for someone to work 20 hours??

OpenStudy (texaschic101):

35 hrs per week, 5 days a week = 35/5 = 7 hrs per day So he can work 2 days and that is 14 hrs. Now he just needs 6 hrs.... and he works 1 1/4 hrs per day... 6/ (1 1/4) = 6/ (5/4) = 6 * 4/5 = 24/5 = 4 4/5 days for a total of : 7 + 4 4/5 = 7 + 24/5 = 35/5 + 24/5 = 59/5 = 11 4/5 days You might want to get a second opinion on this..